OpenStudy (anonymous):

What is the area of the triangle formed by connecting the points (0, 0, 0), (1, 1, 2), and (3, 4, 2)?

OpenStudy (anonymous):

I am thinking heron's formula, but it would be a large number involving many square roots...

OpenStudy (anonymous):

Well I would use the vector product and divide the result by two.

OpenStudy (anonymous):

For instance you can setup two vectors, the origin and the point (1,1,2) and the origin and the point (3,4,2)

OpenStudy (anonymous):

as @waterineyes, you will get the area of a Parallelogram, but that's great because if you divide that by 2 you'll do fine.

OpenStudy (anonymous):

I have not done vectors yet...

OpenStudy (anonymous):

Then you have to go with Herons formula..

OpenStudy (anonymous):

hmm okay, that's one 'easy' way I could suggest, I wouldn't have another idea as of now, because this is 3 Dimensional.

OpenStudy (anonymous):

Alright. Let's see where that leads...

OpenStudy (anonymous):

Firstly calculate the distance of all the sides..

OpenStudy (anonymous):

They are \(\sqrt{6}\), \(\sqrt{13}\), and \(\sqrt{29}\).

OpenStudy (anonymous):

*lengths..

OpenStudy (anonymous):

You have to take it in decimal forms.. Because you are now to find Semi perimter..

OpenStudy (anonymous):

Also no calculators...I'm going to try something else for a second...

OpenStudy (anonymous):

|3 4 2| |1 1 2| |0 0 0| Is there a way to do it with the way you do coordinate geometry?

OpenStudy (anonymous):

I think so but not sure..

OpenStudy (anonymous):

\[\tiny\sqrt{\frac{\sqrt{6} + \sqrt{13} + \sqrt{29}}{2}(\frac{\sqrt{6} + \sqrt{13} + \sqrt{29}}{2} - \sqrt{6})(\frac{\sqrt{6} + \sqrt{13} + \sqrt{29}}{2} - \sqrt{13})(\frac{\sqrt{6} + \sqrt{13} + \sqrt{29}}{2} - \sqrt{29}}\]That's very ugly :/

OpenStudy (anonymous):

Ha ha ha.. Take them in decimals form..

OpenStudy (anonymous):

lol...alright...I'll break the rules until I learn vectors :/

OpenStudy (anonymous):

s ≈ 5.72

OpenStudy (anonymous):

Now use the heron's formula for Area..

OpenStudy (anonymous):

a ≈ 2.45 b ≈ 3.6 c ≈ 5.385 3.27(5.72)(2.12)(0.335) = 13.28?

OpenStudy (anonymous):

I have not checked it if your calculations are right or not..

OpenStudy (anonymous):

Let me check..

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=Area+%280%2C+0%2C+0%29+%281%2C+1%2C+2%29+%283%2C+4%2C+2%29 Wolfram doesn't agree with me...

OpenStudy (anonymous):

There is one square root outside too I guess..

OpenStudy (anonymous):

Oh yeah :/ It works now :)

OpenStudy (anonymous):

THank you! I'm guessing that if I knew vectors it would be a lot easier?

OpenStudy (anonymous):

Just wait.. It will be easy when you will study about vectors.. Welcome dear..